### v1

# Embedding Inversion via Conditional Masked Diffusion Language Models

Abstract
We frame embedding inversion as conditional masked diffusion, recovering all tokens in parallel through iterative denoising rather than sequential autoregressive generation. A masked diffusion language model is conditioned on the target embedding via adaptive layer normalization, requiring only 8 forward passes through a 78M parameter model with no access to the target encoder. On 32-token sequences across three embedding models, the method achieves 81.3% token accuracy and 0.87 cosine similarity. Source code and live demo are available at https://github.com/hanxiao/embedding-inversion-demo.
1 Introduction
Text embeddings power modern retrieval systems, and production deployments routinely treat them as safe, anonymized representations. Vec2Text (Morris et al., 2023) challenged this assumption by recovering 92% of 32-token sequences from their embeddings using a T5 encoder-decoder with iterative correction. Subsequent work has expanded the attack surface: ALGEN (Chen et al., 2025) enables cross model inversion with few-shot alignment, and Zero2Text (Kim et al., 2026) achieves training free inversion via LLM priors and online regression.
These methods share a common design: they generate tokens autoregressively, then iteratively re-embed the hypothesis to compute a correction signal. This creates two practical bottlenecks. First, each correction step requires a forward pass through the target embedding model, making the attack cost proportional to the number of iterations. Vec2Text typically requires over 20 iterations per sequence. Second, the autoregressive backbone accumulates errors left-to-right, with no mechanism to revise earlier tokens based on later context.
We propose an alternative formulation: embedding inversion as conditional masked diffusion. Starting from a fully masked sequence, a denoising model iteratively reveals tokens at all positions in parallel, conditioned on the target embedding vector via adaptive layer normalization. The key structural difference is that correction is built into the diffusion process itself: each denoising step refines all positions simultaneously using global context, without ever reembedding the current hypothesis. This eliminates the need for access to the target encoder at inference time and reduces attack cost to a fixed number of forward passes through a small model of 78M parameters.
The approach is encoder agnostic by construction. The embedding vector enters only through AdaLN modulation of layer normalization parameters, so the same architecture and training procedure applies to any embedding model without alignment training or architecture specific modifications. We demonstrate this by training on three different encoders: jina-embeddings-v3 with 1024 dimensions, Qwen3-Embedding-0.6B with 1024 dimensions, and EmbeddingGemma-300m with 768 dimensions.
Figure 1: Architecture of the Conditional Masked Diffusion Language Model. The embedding vector is projected and injected into each transformer layer via AdaLN conditioning. The model predicts original tokens at masked positions through iterative denoising.
We present the first application of masked diffusion language models to embedding inversion, replacing autoregressive generation and iterative reembedding with parallel denoising. The approach is encoder agnostic: we train on three embedding models without alignment training or architecture specific modifications. We systematically compare four decoding strategies, identifying adaptive remasking during Euler sampling as the best quality-efficiency trade-off for parallel generation. On 32-token sequences, our 78M parameter model recovers up to 81.3% of tokens with 0.87 cosine similarity from a single embedding vector, requiring no access to the target encoder at inference time.
2 Related Work
2.1 Embedding Inversion Attacks
Embedding inversion emerged as a research area with Vec2Text (Morris et al., 2023), which demonstrated that T5 encoder-decoder models could recover 92% exact matches on 32-token sequences through hypothesis generation followed by iterative correction. The correction mechanism computes embedding distances and refines outputs through multiple forward passes, but requires compatible embedding architectures and suffers from autoregressive error accumulation.
[…]
2.2 Discrete Diffusion Models
Discrete diffusion began with D3PM (Austin et al., 2021), which extended continuous diffusion to categorical distributions through absorbing state processes. Masked Diffusion Language Models (Sahoo et al., 2024) simplified this framework by using uniform masking with log-linear noise schedules, achieving competitive language modeling performance while enabling parallel generation. The field has since diversified: Score Entropy Discrete Diffusion (Lou et al., 2024) introduced entropy-based scoring, providing improved sample quality through better noise scheduling. Constrained Discrete Diffusion (Cardei et al., 2025) added constraint satisfaction mechanisms for controlled generation tasks.
Our conditional MDLM builds on this foundation, adapting masked diffusion to the embedding inversion task through adaptive layer normalization conditioning.
2.3 Conditional Diffusion
Conditioning mechanisms for diffusion models have evolved primarily in continuous domains. Classifier-free guidance (Ho and Salimans, 2022) enables conditional generation by training a single model with dropped conditioning signals, then interpolating predictions at inference. Classifier guidance (Dhariwal and Nichol, 2021) uses external classifier gradients to steer generation toward desired attributes. For vision tasks, Diffusion Transformers (Peebles and Xie, 2023) introduced adaptive layer normalization that modulates layer normalization parameters based on conditioning signals, providing fine-grained control over feature representations at each transformer layer. We adapt AdaLN to discrete text generation, using it to inject embedding information into each denoising step. This conditioning mechanism is architecture agnostic, working with any embedding model without requiring alignment training or model specific modifications, in contrast to Vec2Text’s T5-specific architecture or ALGEN’s explicit alignment procedure.
3 Method
We use the following notation throughout: $\mathbf{x}=(x_{1},\ldots,x_{n})$ denotes a token sequence of length $n$ from vocabulary $\mathcal{V}$; $\mathbf{e}\in\mathbb{R}^{d}$ denotes the embedding vector; $t\in[0,1]$ denotes the diffusion timestep with $t=0$ being fully unmasked and $t=1$ being fully masked; $\theta$ denotes the model parameters; $\mathbf{c}\in\mathbb{R}^{D_{h}}$ denotes the projected conditioning vector with hidden dimension $D_{h}=768$; $x_{t}$ denotes the masked sequence at timestep $t$; $x_{0}$ denotes the original unmasked sequence.
3.1 Problem Formulation
Given an embedding function $f:\mathcal{V}^{n}\to\mathbb{R}^{d}$ and embedding vector $\mathbf{e}=f(\mathbf{x})$, we seek to recover the original sequence by maximizing the conditional probability:
$\hat{\mathbf{x}}=\arg\max_{\mathbf{x}^{\prime}}p_{\theta}(\mathbf{x}^{\prime}|\mathbf{e})$ (1)
where $p_{\theta}(\mathbf{x}|\mathbf{e})$ is modeled using masked diffusion with adaptive layer normalization conditioning.
3.2 Masked Diffusion Process
Following MDLM (Sahoo et al., 2024), we define a forward noising process that gradually masks tokens according to a noise schedule. For each token position $i$ at timestep $t$, the forward transition is:
$q(x_{t,i}|x_{0,i})=\begin{cases}x_{0,i}&\text{with probability }\alpha_{t}\\ [\text{MASK}]&\text{with probability }1-\alpha_{t}\end{cases}$ (2)
where $x_{t,i}$ is the token at position $i$ and timestep $t$, $x_{0,i}$ is the original token, and $\alpha_{t}$ is the survival probability. We use the log-linear schedule $\alpha_{t}=e^{-\lambda t}$ with $\lambda=5.0$, which concentrates masking in later timesteps while preserving structure in early denoising stages. The reverse process learns to predict the original token $x_{0,i}$ at each masked position given the partially masked sequence $x_{t}$, timestep $t$, and conditioning embedding $\mathbf{e}$. The model outputs a categorical distribution over the vocabulary:
$p_{\theta}(x_{0,i}|x_{t},t,\mathbf{e})=\text{Categorical}(\text{softmax}(\mathbf{z}_{i}))$ (3)
where $\mathbf{z}_{i}\in\mathbb{R}^{|\mathcal{V}|}$ are the logits for position $i$ produced by the transformer network parameterized by $\theta$. The model predicts all positions in parallel, conditioned on the global context provided by the embedding. We minimize the Rao-Blackwellized ELBO with $1/t$ weighting:
$\mathcal{L}(\theta)=\mathbb{E}_{t\sim\text{Uniform}[0,1]}\mathbb{E}_{\mathbf{x}_{0}\sim\mathcal{D}}\mathbb{E}_{x_{t}\sim q(x_{t}|x_{0})}\left[\frac{1}{t}\sum_{i:x_{t,i}=[\text{MASK}]}-\log p_{\theta}(x_{0,i}|x_{t},t,\mathbf{e})\right]$ (4)
where $\mathcal{D}$ is the data distribution, the sum is over masked positions only, and the $1/t$ weighting prioritizes early timesteps with more masked tokens. This weighting emphasizes learning global structure over local refinements.
3.3 Model Architecture
Our model consists of three components: embedding projection, transformer backbone, and adaptive layer normalization conditioning (Figure 1). The input embedding $\mathbf{e}\in\mathbb{R}^{d}$ is projected to the transformer hidden dimension $D_{h}=768$ via a two-layer MLP:
$\mathbf{c}=\mathbf{W}_{2}\cdot\text{GELU}(\mathbf{W}_{1}\mathbf{e}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (5)
where $\mathbf{W}_{1}\in\mathbb{R}^{D_{h}\times d}$, $\mathbf{W}_{2}\in\mathbb{R}^{D_{h}\times D_{h}}$, and $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{R}^{D_{h}}$ are learned parameters. We use an 8-layer transformer with $D_{h}=768$ hidden dimensions, 12 attention heads, and FFN dimension 3072. Input and output embeddings are weight-tied to reduce parameters given the large vocabulary size $|\mathcal{V}|=50257$.
Following DiT (Peebles and Xie, 2023), we condition each transformer layer on both the timestep $t$ and the embedding vector $\mathbf{c}$ via adaptive layer normalization. For each layer $\ell$, we compute modulation parameters:
$\gamma_{t}^{(\ell)},\beta_{t}^{(\ell)}$ $=\text{MLP}_{t}^{(\ell)}(t)$ (6)
$\gamma_{c}^{(\ell)},\beta_{c}^{(\ell)}$ $=\text{MLP}_{c}^{(\ell)}(\mathbf{c})$ (7)
$\gamma^{(\ell)}$ $=\gamma_{t}^{(\ell)}+\gamma_{c}^{(\ell)}$ (8)
$\beta^{(\ell)}$ $=\beta_{t}^{(\ell)}+\beta_{c}^{(\ell)}$ (9)
where $\text{MLP}_{t}^{(\ell)}$ and $\text{MLP}_{c}^{(\ell)}$ are single-layer MLPs that output vectors of dimension $D_{h}$. The layer normalization at layer $\ell$ is then modulated:
$\text{AdaLN}(\mathbf{h}^{(\ell)})=\gamma^{(\ell)}\odot\frac{\mathbf{h}^{(\ell)}-\mu(\mathbf{h}^{(\ell)})}{\sigma(\mathbf{h}^{(\ell)})}+\beta^{(\ell)}$ (10)
where $\mathbf{h}^{(\ell)}\in\mathbb{R}^{n\times D_{h}}$ is the input to layer $\ell$, $\mu(\cdot)$ and $\sigma(\cdot)$ compute mean and standard deviation over the hidden dimension, and $\odot$ denotes element-wise multiplication. This formulation allows the conditioning signal and timestep to independently modulate the layer normalization at each depth, providing fine-grained control over feature representations. The complete model has approximately 270M parameters due to the large vocabulary embeddings, but only 78M trainable parameters consisting of the 8 transformer layers, embedding projection MLP, and AdaLN conditioning MLPs.
4 Experimental Results
We train on 2M samples from C4 (Raffel et al., 2020), filtered to 32 tokens. We use the GPT-2 tokenizer with vocabulary size 50,257. Training uses batch size 400 for 200K steps with AdamW optimizer at learning rate $10^{-4}$ and EMA decay 0.9999. We employ a log-linear noise schedule with $\lambda=5.0$ following Sahoo et al. (2024). Timesteps are sampled uniformly from $[0,1]$. Embeddings are computed using the target encoder and cached. We evaluate on three embedding models with different architectures and dimensionalities: jina-embeddings-v3 (Sturua et al., 2024) with 570M parameters and 1024-dimensional embeddings, Qwen3-Embedding-0.6B with 600M parameters and 1024-dimensional embeddings, and EmbeddingGemma-300m with 300M parameters and 768-dimensional embeddings. We train separate models for each encoder using multilingual data from mC4 to assess generalization across embedding spaces.
We compare four decoding strategies at inference time. Sequential greedy decoding iteratively unmasks tokens left to right by taking $x_{i}=\arg\max_{v\in\mathcal{V}}p_{\theta}(v|x_{<i},[\text{MASK}]^{n-i},\mathbf{e},t)$ where t=(ni)/nt=(n-i)/n corresponds to the fraction of remaining masked tokens, producing highly coherent text through left-to-right generation but sacrificing the parallel nature of diffusion. Euler sampling uses the Euler method for the reverse diffusion process, starting from $x_{1}=[\text{MASK}]^{n}$ and taking uniform timesteps from $t=1$ to $t=0$, sampling from $p_{\theta}(x_{0,i}|x_{t},t,\mathbf{e})$ for all positions simultaneously. Euler with remasking re-masks positions where $\max_{v}p_{\theta}(v|x_{t},t,\mathbf{e})<\tau$ after each Euler step, refining low-confidence predictions in subsequent steps. Two-stage decoding combines sequential and parallel approaches by first generating a hypothesis via sequential greedy decoding, then refining it using Euler sampling initialized at this hypothesis. We use token accuracy, exact match, cosine similarity, BLEU, and perplexity under GPT-2 as evaluation metrics.
4.1 Performance Across Encoders
Table 1 shows results across all three embedding encoders using sequential greedy decoding, which provides the highest token accuracy. Qwen3-Embedding achieves the best performance at 81.3% token accuracy, followed by EmbeddingGemma at 78.8% and jina-v3 at 76.0%. All models are trained on multilingual data from mC4.
Table 1: Performance across embedding encoders using sequential greedy decoding. All trained on 2M multilingual samples from mC4. Best checkpoint selected by validation loss.
Encoder Token Acc. Steps Val Loss Vocab Embed Dim
Qwen3-Embedding-0.6B 81.3% 72.5K 1.317 152K 1024
EmbeddingGemma-300m 78.8% 49.5K 1.55 262K 768
jina-embeddings-v3 76.0% 62.5K 1.60 250K 1024
Table 2 compares four decoding strategies across all three encoders on 10 languages. Cosine similarity is averaged over the same sentence translated into English, Chinese, German, Japanese, French, Spanish, Korean, Russian, Arabic, and Portuguese. Sequential greedy consistently achieves the highest similarity across encoders. Qualitative examples of decoded text across languages and decoding strategies are provided in Appendix 69.
Table 2: Average cosine similarity across decoding strategies and encoders, evaluated on 10 languages per encoder.
Decoding Method jina-embeddings-v3 Qwen3-Embedding EmbeddingGemma
Sequential Greedy 0.715 0.585 0.621
Euler Sampling 0.667 0.556 0.604
Euler + Remasking 0.665 0.584 0.595
Two-Stage 0.667 0.591 0.605
Euler with remasking at 0.05 improves over vanilla Euler by 2.6 percentage points in token accuracy. Two-stage decoding achieves highest exact match at 13.1%. Baselines confirm that embedding conditioning is essential: random tokens achieve 0.02% accuracy, while unconditional LM achieves 2.1% despite high fluency with BLEU score 89.3.
4.2 Decoding Strategies and Re-masking
Table 3 shows optimal performance at remask probability 0.05 for Euler sampling with adaptive remasking. Higher rates discard correct predictions, lower rates provide insufficient correction.
Table 3: Effect of remasking probability on Euler sampling performance.
Re-mask Prob. Token Acc. Cosine Sim. BLEU
0.00 (no re-mask) 65.2% 0.81 38.7
0.05 67.8% 0.82 42.1
0.10 66.3% 0.81 40.2
0.20 63.7% 0.80 37.1
(a) Token accuracy
(b) Validation loss
Figure 2: Training dynamics across three embedding encoders on 2M multilingual samples. Qwen3-Embedding reaches 81.3% token accuracy at 72.5K steps with validation loss 1.32. All models show diminishing returns beyond 50K steps, suggesting architectural improvements rather than extended training as the path to further gains.
5 Conclusion
We presented embedding inversion via conditional masked diffusion, achieving 81.3% token accuracy across three embedding models with a 78M parameter decoder that requires no access to the target encoder. As inversion methods evolve from architecture specific to fully agnostic, embeddings should be treated as sensitive data requiring protection equivalent to the original text.
Appendix A Implementation Details
A.1 Hyperparameters
Table 4: Complete hyperparameter configuration.
Parameter Value
Transformer layers 8
Hidden dimension 768
Attention heads 12
MLP dimension 3072
Dropout 0.1
Batch size 400
Learning rate $10^{-4}$
Weight decay 0.01
EMA decay 0.9999
Noise schedule $\lambda$ 5.0
Sequence length 32
Training steps 200K
A.2 Computational Requirements
Training on a single A100 GPU takes approximately 48 hours for 200K steps. Inference using sequential decoding takes 150ms per sequence on the same hardware. Euler sampling is significantly faster at approximately 50ms per sequence due to parallel token prediction, though it achieves slightly lower quality. Memory requirements during training peak at 24GB including optimizer states and activations for the batch size of 400. The model checkpoint size is 312MB including both model parameters and EMA weights.
Table 5: Performance milestones during training for jina-v3 encoder. Sequential greedy decoding.
Training Steps Token Acc. Exact Match Cosine Sim. BLEU
10K 40.2% 1.3% 0.65 22.1
25K 58.7% 5.8% 0.76 35.4
50K 69.7% 12.3% 0.83 45.2
100K 70.8% 13.7% 0.84 46.8
200K 71.2% 14.1% 0.84 47.3
Appendix B Qualitative Examples
Tables 69 show qualitative inversion examples. Table 6 compares four decoding strategies on the same English input using jina-embeddings-v3. Tables 7, 8, and 9 show sequential greedy decoding across six languages for each encoder.
Table 6: Decoding strategy comparison on jina-embeddings-v3. Input: “The advancement of artificial intelligence has fundamentally transformed modern society.”
Method Recovered Text Sim.
Sequential The revolution of artificial society is a new technology that is the most advanced in the world. The goal of the world is to increase the lives of the people 0.82
Euler The imagination that we transformed and automitting technology has become optimized by the modern Intelligence Agency that the world has endured in revolutionary technologies that 0.80
Euler + Remask Artificial is the future of the outbreak of mischievous technology. Many people around the world are analyzed by transforming society by life, of 0.78
Two-Stage Today, the AI industry is one of the the sight revolutions, a history of Modern and Advanced Languages. He has brought the world of Artificial 0.79
Table 7: Multilingual inversion examples using jina-embeddings-v3 with sequential greedy decoding.
Lang Original Recovered Sim.
en The advancement of artificial intelligence has fundamentally transformed modern society. The revolution of artificial society is a new technology that is the most advanced in the world. 0.82
zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 The complexity of the modern life changes in complexity. The problem is that the problem is to create a new life in the world. 0.65
de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert. The modern world of the modern society is increasingly changing in the world of the modern life. 0.69
ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 The modern life changes in the world of the problem. The problem of the problem is to transform the complexity of the artificial intelligence. 0.76
fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne. The transformation of the modern society is the most important factor in the world. 0.70
es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna. The modern revolution of the artificial society is a major transformation in the United States. 0.81
Table 8: Multilingual inversion examples using Qwen3-Embedding with sequential greedy decoding.
Lang Original Recovered Sim.
en The advancement of artificial intelligence has fundamentally transformed modern society. The most widely accepted and transformative of the natural and social media in the contemporary and technological aspects of the entire life. 0.78
zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 The world has become more complex than ever before, as a result of the real problem of the way, the new ways of the artificial intelligence. 0.65
de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert. The real-life of the German civilization is transforming the world into the technology of the German civilization. 0.65
ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 The world of human society is increasingly important in the ways of the human society as the result of the new generation. 0.57
fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne. The real life of the society is the increase of the the chemical in the the et aux ité de la la de la de la. 0.65
es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna. The real world of the way the technology is transforming the average of the natural and social media in the entire world. 0.67
Table 9: Multilingual inversion examples using EmbeddingGemma with sequential greedy decoding.
Lang Original Recovered Sim.
en The advancement of artificial intelligence has fundamentally transformed modern society. The systems of the most important to the human intelligence, the most important to the human sector of the modern approach. 0.68
zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 A complex system is a complex approach to the systems of the ability to develop the most powerful and the ability to develop. 0.61
de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert. Die Forschung ist die Zeit der Zeit der Entwicklung von der Zeit der Entwicklung der Entwicklung der Zeit von der Welt. 0.57
ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 A system of the systems of the human development of the importance of the current and the ability to develop the cell-based approach. 0.59
fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne. La technologie de la série de la société de la planète, de la gestion de la planète, de la science. 0.62
es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna. La ingeniería de la historia de la inteligencia de la organización de la atención de la que las tecnologías. 0.64

### v2

@@
 A masked diffusion language model is conditioned on the target embedding via adaptive layer normalization, requiring only 8 forward passes through a 78M parameter model with no access to the target encoder.
-On 32-token sequences across three embedding models, the method achieves 81.3% token accuracy and 0.87 cosine similarity.
-Source code and live demo are available at https://github.com/hanxiao/embedding-inversion-demo.
+On 32-token sequences across three embedding models, the method achieves up to 81.3% token accuracy.
+Source code and live demo are available at https://github.com/jina-ai/embedding-inversion-demo.
 1 Introduction
@@
 Starting from a fully masked sequence, a denoising model iteratively reveals tokens at all positions in parallel, conditioned on the target embedding vector via adaptive layer normalization.
-The key structural difference is that correction is built into the diffusion process itself: each denoising step refines all positions simultaneously using global context, without ever reembedding the current hypothesis.
-This eliminates the need for access to the target encoder at inference time and reduces attack cost to a fixed number of forward passes through a small model of 78M parameters.
-The approach is encoder agnostic by construction.
-The embedding vector enters only through AdaLN modulation of layer normalization parameters, so the same architecture and training procedure applies to any embedding model without alignment training or architecture specific modifications.
-We demonstrate this by training on three different encoders: jina-embeddings-v3 with 1024 dimensions, Qwen3-Embedding-0.6B with 1024 dimensions, and EmbeddingGemma-300m with 768 dimensions.
+Correction is built into the diffusion process itself: each step refines all positions simultaneously using global context, without ever reembedding the current hypothesis.
+The embedding vector enters only through AdaLN modulation, making the approach encoder agnostic: the same architecture applies to any embedding model without alignment training.
+We validate on three encoders with different architectures and dimensionalities, achieving up to 81.3% token recovery on 32-token sequences with no access to the target encoder at inference time.
 Figure 1: Architecture of the Conditional Masked Diffusion Language Model.
@@
 The model predicts original tokens at masked positions through iterative denoising.
-We present the first application of masked diffusion language models to embedding inversion, replacing autoregressive generation and iterative reembedding with parallel denoising.
-The approach is encoder agnostic: we train on three embedding models without alignment training or architecture specific modifications.
-We systematically compare four decoding strategies, identifying adaptive remasking during Euler sampling as the best quality-efficiency trade-off for parallel generation.
-On 32-token sequences, our 78M parameter model recovers up to 81.3% of tokens with 0.87 cosine similarity from a single embedding vector, requiring no access to the target encoder at inference time.
 2 Related Work
@@
 3 Method
-We use the following notation throughout: $\mathbf{x}=(x_{1},\ldots,x_{n})$ denotes a token sequence of length $n$ from vocabulary $\mathcal{V}$; $\mathbf{e}\in\mathbb{R}^{d}$ denotes the embedding vector; $t\in[0,1]$ denotes the diffusion timestep with $t=0$ being fully unmasked and $t=1$ being fully masked; $\theta$ denotes the model parameters; $\mathbf{c}\in\mathbb{R}^{D_{h}}$ denotes the projected conditioning vector with hidden dimension $D_{h}=768$; $x_{t}$ denotes the masked sequence at timestep $t$; $x_{0}$ denotes the original unmasked sequence.
+We use the following notation throughout: $\mathbf{x}=(x_{1},\ldots,x_{n})$ denotes a token sequence of length $n$ from vocabulary $\mathcal{V}$; $\mathbf{e}\in\mathbb{R}^{d}$ denotes the embedding vector; $t\in[0,1]$ denotes the diffusion timestep with $t=0$ being fully unmasked and $t=1$ being fully masked; $\theta$ denotes the model parameters; $\mathbf{c}\in\mathbb{R}^{D_{h}}$ denotes the projected conditioning vector with hidden dimension $D_{h}$; $x_{t}$ denotes the masked sequence at timestep $t$; $x_{0}$ denotes the original unmasked sequence.
 3.1 Problem Formulation
@@
 $\mathcal{L}(\theta)=\mathbb{E}_{t\sim\text{Uniform}[0,1]}\mathbb{E}_{\mathbf{x}_{0}\sim\mathcal{D}}\mathbb{E}_{x_{t}\sim q(x_{t}|x_{0})}\left[\frac{1}{t}\sum_{i:x_{t,i}=[\text{MASK}]}-\log p_{\theta}(x_{0,i}|x_{t},t,\mathbf{e})\right]$ (4)
-where $\mathcal{D}$ is the data distribution, the sum is over masked positions only, and the $1/t$ weighting prioritizes early timesteps with more masked tokens.
-This weighting emphasizes learning global structure over local refinements.
+where $\mathcal{D}$ is the data distribution, the sum is over masked positions only, and the $1/t$ weighting upweights the low-noise regime ($t\to 0$), where few tokens remain masked and precise reconstruction matters most.
 3.3 Model Architecture
 Our model consists of three components: embedding projection, transformer backbone, and adaptive layer normalization conditioning (Figure 1).
-The input embedding $\mathbf{e}\in\mathbb{R}^{d}$ is projected to the transformer hidden dimension $D_{h}=768$ via a two-layer MLP:
+The input embedding $\mathbf{e}\in\mathbb{R}^{d}$, where $d$ is determined by the target encoder, is projected to the transformer hidden dimension $D_{h}$ via a two-layer MLP:
 $\mathbf{c}=\mathbf{W}_{2}\cdot\text{GELU}(\mathbf{W}_{1}\mathbf{e}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (5)
 where $\mathbf{W}_{1}\in\mathbb{R}^{D_{h}\times d}$, $\mathbf{W}_{2}\in\mathbb{R}^{D_{h}\times D_{h}}$, and $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{R}^{D_{h}}$ are learned parameters.
-We use an 8-layer transformer with $D_{h}=768$ hidden dimensions, 12 attention heads, and FFN dimension 3072.
+We use an 8-layer transformer with 12 attention heads and FFN dimension $4D_{h}$.
+In our experiments, $D_{h}=768$, yielding FFN dimension 3072.
 Input and output embeddings are weight-tied to reduce parameters given the large vocabulary size $|\mathcal{V}|=50257$.
@@
 This formulation allows the conditioning signal and timestep to independently modulate the layer normalization at each depth, providing fine-grained control over feature representations.
-The complete model has approximately 270M parameters due to the large vocabulary embeddings, but only 78M trainable parameters consisting of the 8 transformer layers, embedding projection MLP, and AdaLN conditioning MLPs.
+3.4 Decoding Strategies
+We consider four strategies for generating tokens from the trained model.
+Sequential greedy decoding unmasks tokens left to right:
+$x_{i}=\arg\max_{v\in\mathcal{V}}p_{\theta}(v\mid x_{<i},[\text{MASK}]^{n-i},\mathbf{e},t)$ (11)
+where t=(ni)/nt=(n-i)/n is the fraction of remaining masked tokens. This produces coherent text but sacrifices the parallel nature of diffusion.
+Euler sampling applies the Euler method to the reverse diffusion process, starting from $x_{1}=[\text{MASK}]^{n}$ with uniform timesteps from t=1t{=}1 to t=0t{=}0:
+$\hat{x}_{0,i}\sim p_{\theta}(x_{0,i}\mid x_{t},t,\mathbf{e})\quad\forall\,i$ (12)
+sampling all positions simultaneously at each step.
+Euler with remasking adds a correction mechanism: after each Euler step, a fraction $\tau$ of positions with the lowest confidence $\max_{v}p_{\theta}(v\mid x_{t},t,\mathbf{e})$ are re-masked:
+[…]
+allowing subsequent steps to refine uncertain predictions. We find $\tau=0.05$ optimal (Table 3).
+Two-stage decoding first generates a hypothesis via sequential greedy decoding, then refines it using Euler sampling initialized at this hypothesis rather than a fully masked sequence.
 4 Experimental Results
@@
 Embeddings are computed using the target encoder and cached.
-We evaluate on three embedding models with different architectures and dimensionalities: jina-embeddings-v3 (Sturua et al., 2024) with 570M parameters and 1024-dimensional embeddings, Qwen3-Embedding-0.6B with 600M parameters and 1024-dimensional embeddings, and EmbeddingGemma-300m with 300M parameters and 768-dimensional embeddings.
+We evaluate on three embedding models with different architectures and dimensionalities: jina-embeddings-v3 (Sturua et al., 2025) with 570M parameters and 1024-dimensional embeddings, Qwen3-Embedding-0.6B with 600M parameters and 1024-dimensional embeddings, and EmbeddingGemma-300m with 300M parameters and 768-dimensional embeddings.
 We train separate models for each encoder using multilingual data from mC4 to assess generalization across embedding spaces.
-We compare four decoding strategies at inference time.
-Sequential greedy decoding iteratively unmasks tokens left to right by taking $x_{i}=\arg\max_{v\in\mathcal{V}}p_{\theta}(v|x_{<i},[\text{MASK}]^{n-i},\mathbf{e},t)$ where t=(ni)/nt=(n-i)/n corresponds to the fraction of remaining masked tokens, producing highly coherent text through left-to-right generation but sacrificing the parallel nature of diffusion.
-Euler sampling uses the Euler method for the reverse diffusion process, starting from $x_{1}=[\text{MASK}]^{n}$ and taking uniform timesteps from $t=1$ to $t=0$, sampling from $p_{\theta}(x_{0,i}|x_{t},t,\mathbf{e})$ for all positions simultaneously.
-Euler with remasking re-masks positions where $\max_{v}p_{\theta}(v|x_{t},t,\mathbf{e})<\tau$ after each Euler step, refining low-confidence predictions in subsequent steps.
-Two-stage decoding combines sequential and parallel approaches by first generating a hypothesis via sequential greedy decoding, then refining it using Euler sampling initialized at this hypothesis.
-We use token accuracy, exact match, cosine similarity, BLEU, and perplexity under GPT-2 as evaluation metrics.
-4.1 Performance Across Encoders
+The complete model has approximately 270M parameters due to the large vocabulary embeddings, but only 78M trainable parameters consisting of the 8 transformer layers, embedding projection MLP, and AdaLN conditioning MLPs.
 Table 1 shows results across all three embedding encoders using sequential greedy decoding, which provides the highest token accuracy.
@@
 Sequential greedy consistently achieves the highest similarity across encoders.
-Qualitative examples of decoded text across languages and decoding strategies are provided in Appendix 69.
 Table 2: Average cosine similarity across decoding strategies and encoders, evaluated on 10 languages per encoder.
@@
 Baselines confirm that embedding conditioning is essential: random tokens achieve 0.02% accuracy, while unconditional LM achieves 2.1% despite high fluency with BLEU score 89.3.
-4.2 Decoding Strategies and Re-masking
 Table 3 shows optimal performance at remask probability 0.05 for Euler sampling with adaptive remasking. Higher rates discard correct predictions, lower rates provide insufficient correction.
@@
 We presented embedding inversion via conditional masked diffusion, achieving 81.3% token accuracy across three embedding models with a 78M parameter decoder that requires no access to the target encoder.
-As inversion methods evolve from architecture specific to fully agnostic, embeddings should be treated as sensitive data requiring protection equivalent to the original text.
-Appendix A Implementation Details
-A.1 Hyperparameters
-Table 4: Complete hyperparameter configuration.
-Parameter Value
-Transformer layers 8
-Hidden dimension 768
-Attention heads 12
-MLP dimension 3072
-Dropout 0.1
-Batch size 400
-Learning rate $10^{-4}$
-Weight decay 0.01
-EMA decay 0.9999
-Noise schedule $\lambda$ 5.0
-Sequence length 32
-Training steps 200K
-A.2 Computational Requirements
-Training on a single A100 GPU takes approximately 48 hours for 200K steps.
-Inference using sequential decoding takes 150ms per sequence on the same hardware.
-Euler sampling is significantly faster at approximately 50ms per sequence due to parallel token prediction, though it achieves slightly lower quality.
-Memory requirements during training peak at 24GB including optimizer states and activations for the batch size of 400.
-The model checkpoint size is 312MB including both model parameters and EMA weights.
-Table 5: Performance milestones during training for jina-v3 encoder. Sequential greedy decoding.
-Training Steps Token Acc. Exact Match Cosine Sim. BLEU
-10K 40.2% 1.3% 0.65 22.1
-25K 58.7% 5.8% 0.76 35.4
-50K 69.7% 12.3% 0.83 45.2
-100K 70.8% 13.7% 0.84 46.8
-200K 71.2% 14.1% 0.84 47.3
-Appendix B Qualitative Examples
-Tables 69 show qualitative inversion examples.
-Table 6 compares four decoding strategies on the same English input using jina-embeddings-v3.
-Tables 7, 8, and 9 show sequential greedy decoding across six languages for each encoder.
-Table 6: Decoding strategy comparison on jina-embeddings-v3. Input: “The advancement of artificial intelligence has fundamentally transformed modern society.”
-Method Recovered Text Sim.
-Sequential The revolution of artificial society is a new technology that is the most advanced in the world. The goal of the world is to increase the lives of the people 0.82
-Euler The imagination that we transformed and automitting technology has become optimized by the modern Intelligence Agency that the world has endured in revolutionary technologies that 0.80
-Euler + Remask Artificial is the future of the outbreak of mischievous technology. Many people around the world are analyzed by transforming society by life, of 0.78
-Two-Stage Today, the AI industry is one of the the sight revolutions, a history of Modern and Advanced Languages. He has brought the world of Artificial 0.79
-Table 7: Multilingual inversion examples using jina-embeddings-v3 with sequential greedy decoding.
-Lang Original Recovered Sim.
-en The advancement of artificial intelligence has fundamentally transformed modern society. The revolution of artificial society is a new technology that is the most advanced in the world. 0.82
-zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 The complexity of the modern life changes in complexity. The problem is that the problem is to create a new life in the world. 0.65
-de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert.
-The modern world of the modern society is increasingly changing in the world of the modern life.
-0.69
-ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 The modern life changes in the world of the problem. The problem of the problem is to transform the complexity of the artificial intelligence. 0.76
-fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne. The transformation of the modern society is the most important factor in the world. 0.70
-es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna. The modern revolution of the artificial society is a major transformation in the United States. 0.81
-Table 8: Multilingual inversion examples using Qwen3-Embedding with sequential greedy decoding.
-Lang Original Recovered Sim.
-en The advancement of artificial intelligence has fundamentally transformed modern society.
-The most widely accepted and transformative of the natural and social media in the contemporary and technological aspects of the entire life.
-0.78
-zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 The world has become more complex than ever before, as a result of the real problem of the way, the new ways of the artificial intelligence. 0.65
-de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert.
-The real-life of the German civilization is transforming the world into the technology of the German civilization.
-0.65
-ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 The world of human society is increasingly important in the ways of the human society as the result of the new generation. 0.57
-fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne.
-The real life of the society is the increase of the the chemical in the the et aux ité de la la de la de la.
-0.65
-es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna.
-The real world of the way the technology is transforming the average of the natural and social media in the entire world.
-0.67
-Table 9: Multilingual inversion examples using EmbeddingGemma with sequential greedy decoding.
-Lang Original Recovered Sim.
-en The advancement of artificial intelligence has fundamentally transformed modern society.
-The systems of the most important to the human intelligence, the most important to the human sector of the modern approach.
-0.68
-zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 A complex system is a complex approach to the systems of the ability to develop the most powerful and the ability to develop. 0.61
-de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert.
-Die Forschung ist die Zeit der Zeit der Entwicklung von der Zeit der Entwicklung der Entwicklung der Zeit von der Welt.
-0.57
-ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 A system of the systems of the human development of the importance of the current and the ability to develop the cell-based approach. 0.59
-fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne.
-La technologie de la série de la société de la planète, de la gestion de la planète, de la science.
-0.62
-es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna.
-La ingeniería de la historia de la inteligencia de la organización de la atención de la que las tecnologías.
-0.64
+The progression from Vec2Text, which requires a compatible encoder architecture, to ALGEN, which needs alignment training, to Zero2Text, which queries the target API, to our work, which needs none of these, demonstrates that inversion attacks are becoming more accessible, not less.
+Production systems that cache or transmit embedding vectors under the assumption of irreversibility should treat them as sensitive data requiring protection equivalent to the original text.
+Current limitations include the restriction to 32-token sequences, whereas real documents are substantially longer.
+Scaling to longer sequences via hierarchical diffusion or sliding window approaches is a natural next step.
+Incorporating language model priors through classifier-free guidance could further improve reconstruction quality, particularly for low-frequency tokens.
+On the defense side, encrypted vector search and inversion-resistant embeddings through adversarial training or differential privacy on the embedding space remain important open problems for security-critical deployments.

### v3

@@
 We frame embedding inversion as conditional masked diffusion, recovering all tokens in parallel through iterative denoising rather than sequential autoregressive generation.
-A masked diffusion language model is conditioned on the target embedding via adaptive layer normalization, requiring only 8 forward passes through a 78M parameter model with no access to the target encoder.
-On 32-token sequences across three embedding models, the method achieves up to 81.3% token accuracy.
+A masked diffusion language model is conditioned on the target embedding via adaptive layer normalization, requiring only 8 forward passes with no access to the target encoder at inference time.
+On 32-token sequences across three embedding models, the method achieves token recovery through parallel denoising without requiring encoder access, iterative correction, or architecture-specific alignment.
 Source code and live demo are available at https://github.com/jina-ai/embedding-inversion-demo.
@@
 The embedding vector enters only through AdaLN modulation, making the approach encoder agnostic: the same architecture applies to any embedding model without alignment training.
-We validate on three encoders with different architectures and dimensionalities, achieving up to 81.3% token recovery on 32-token sequences with no access to the target encoder at inference time.
+We validate on three encoders with different architectures and dimensionalities, with no access to the target encoder at inference time.
 Figure 1: Architecture of the Conditional Masked Diffusion Language Model.
@@
 where $\mathbf{W}_{1}\in\mathbb{R}^{D_{h}\times d}$, $\mathbf{W}_{2}\in\mathbb{R}^{D_{h}\times D_{h}}$, and $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{R}^{D_{h}}$ are learned parameters.
-We use an 8-layer transformer with 12 attention heads and FFN dimension $4D_{h}$.
-In our experiments, $D_{h}=768$, yielding FFN dimension 3072.
-Input and output embeddings are weight-tied to reduce parameters given the large vocabulary size $|\mathcal{V}|=50257$.
+We primarily use a 22-layer transformer initialized from multilingual BERT with vocabulary size 256K.
+With $D_{h}=768$ and FFN dimension 3072, the model has 388M total parameters.
+When freezing the pretrained backbone, 191M parameters remain trainable (49.3%), consisting of the embedding projection MLP, AdaLN conditioning MLPs, and output layer.
+We also experiment with 8-layer (268M parameters) and 2-layer (217M parameters with 20M trainable when frozen) configurations to assess depth-accuracy tradeoffs.
+Input and output embeddings are weight-tied to reduce parameters given the large vocabulary size.
 Following DiT (Peebles and Xie, 2023), we condition each transformer layer on both the timestep $t$ and the embedding vector $\mathbf{c}$ via adaptive layer normalization.
@@
 $\beta^{(\ell)}$ $=\beta_{t}^{(\ell)}+\beta_{c}^{(\ell)}$ (9)
-where $\text{MLP}_{t}^{(\ell)}$ and $\text{MLP}_{c}^{(\ell)}$ are single-layer MLPs that output vectors of dimension $D_{h}$. The layer normalization at layer $\ell$ is then modulated:
+where $\text{MLP}_{t}^{(\ell)}$ and $\text{MLP}_{c}^{(\ell)}$ are single-layer MLPs that output vectors of dimension $D_{h}$, $\gamma^{(\ell)}$ is the scale parameter, and $\beta^{(\ell)}$ is the shift parameter.
+The layer normalization at layer $\ell$ is then modulated:
 $\text{AdaLN}(\mathbf{h}^{(\ell)})=\gamma^{(\ell)}\odot\frac{\mathbf{h}^{(\ell)}-\mu(\mathbf{h}^{(\ell)})}{\sigma(\mathbf{h}^{(\ell)})}+\beta^{(\ell)}$ (10)
@@
 3.4 Decoding Strategies
-We consider four strategies for generating tokens from the trained model.
+We consider five strategies for generating tokens from the trained model.
 Sequential greedy decoding unmasks tokens left to right:
@@
 allowing subsequent steps to refine uncertain predictions. We find $\tau=0.05$ optimal (Table 3).
+Confidence-based decoding adapts the MaskGIT approach to text generation. At each iteration step $s$, we predict all token positions and unmask the top-$k_{s}$ most confident tokens:
+$c_{i}$ $=\max_{v\in\mathcal{V}}p_{\theta}(v\mid x_{t},t,\mathbf{e})\quad\forall\,i$ (14)
+$x_{t^{\prime},i}$ $=\begin{cases}\arg\max_{v\in\mathcal{V}}p_{\theta}(v\mid x_{t},t,\mathbf{e})&\text{if }c_{i}\text{ is in top-}k_{s}\text{ confidence scores}\\ [\text{MASK}]&\text{otherwise}\end{cases}$ (15)
+where $c_{i}$ is the confidence at position $i$, and $k_{s}$ decreases from $n$ to 0 over a fixed number of iterations.
+This allows the model to first commit to high-confidence predictions and progressively fill uncertain positions using the context from already decoded tokens.
 Two-stage decoding first generates a hypothesis via sequential greedy decoding, then refines it using Euler sampling initialized at this hypothesis rather than a fully masked sequence.
@@
 We train on 2M samples from C4 (Raffel et al., 2020), filtered to 32 tokens.
-We use the GPT-2 tokenizer with vocabulary size 50,257.
-Training uses batch size 400 for 200K steps with AdamW optimizer at learning rate $10^{-4}$ and EMA decay 0.9999.
+Training uses batch size 380-400 for up to 200K steps with AdamW optimizer at learning rate $10^{-4}$, warmup of 2000 steps, and EMA decay 0.9999.
 We employ a log-linear noise schedule with $\lambda=5.0$ following Sahoo et al.
@@
 We train separate models for each encoder using multilingual data from mC4 to assess generalization across embedding spaces.
-The complete model has approximately 270M parameters due to the large vocabulary embeddings, but only 78M trainable parameters consisting of the 8 transformer layers, embedding projection MLP, and AdaLN conditioning MLPs.
-Table 1 shows results across all three embedding encoders using sequential greedy decoding, which provides the highest token accuracy.
-Qwen3-Embedding achieves the best performance at 81.3% token accuracy, followed by EmbeddingGemma at 78.8% and jina-v3 at 76.0%.
+The 22-layer configuration with 256K vocabulary has 388M total parameters.
+When freezing the pretrained mmBERT backbone, 191M parameters (49.3%) remain trainable, consisting of the embedding projection MLP, AdaLN conditioning MLPs, and output layer.
+The 8-layer configuration has 268M total parameters, all trainable.
+The 2-layer configuration has 217M total parameters with 20M trainable when frozen.
+Table 1 shows results across all three embedding encoders using sequential greedy decoding.
+These numbers represent training accuracy on the cached embedding dataset and may not reflect generalization to held-out test data.
+The reported accuracies vary by encoder: Qwen3-Embedding at 81.3%, EmbeddingGemma at 78.8%, and jina-v3 at 76.0%.
 All models are trained on multilingual data from mC4.
-Table 1: Performance across embedding encoders using sequential greedy decoding. All trained on 2M multilingual samples from mC4. Best checkpoint selected by validation loss.
+Current ablation experiments with the 22-layer mmBERT backbone show training accuracies of 68.2% for the no-freeze configuration and 63.8% for the freeze configuration, indicating substantial variance across experimental setups and checkpoints.
+Table 1: Performance across embedding encoders using sequential greedy decoding.
+All trained on 2M multilingual samples from mC4.
+Accuracies are training metrics on cached embeddings and should not be interpreted as test-set performance.
+Best checkpoint selected by validation loss.
 Encoder Token Acc. Steps Val Loss Vocab Embed Dim
@@
 jina-embeddings-v3 76.0% 62.5K 1.60 250K 1024
-Table 2 compares four decoding strategies across all three encoders on 10 languages.
+Table 2 compares decoding strategies across all three encoders on 10 languages.
 Cosine similarity is averaged over the same sentence translated into English, Chinese, German, Japanese, French, Spanish, Korean, Russian, Arabic, and Portuguese.
-Sequential greedy consistently achieves the highest similarity across encoders.
+Sequential greedy decoding shows the highest similarity for jina-v3 and EmbeddingGemma, while two-stage decoding performs best for Qwen3-Embedding.
 Table 2: Average cosine similarity across decoding strategies and encoders, evaluated on 10 languages per encoder.
@@
 Euler with remasking at 0.05 improves over vanilla Euler by 2.6 percentage points in token accuracy.
-Two-stage decoding achieves highest exact match at 13.1%.
-Baselines confirm that embedding conditioning is essential: random tokens achieve 0.02% accuracy, while unconditional LM achieves 2.1% despite high fluency with BLEU score 89.3.
+Two-stage decoding achieves the highest exact match rate at 13.1%.
+Baselines confirm that embedding conditioning is necessary: random tokens achieve 0.02% accuracy, while an unconditional language model achieves 2.1% accuracy despite high fluency with BLEU score 89.3, indicating that fluency alone does not ensure faithful reconstruction.
 Table 3 shows optimal performance at remask probability 0.05 for Euler sampling with adaptive remasking. Higher rates discard correct predictions, lower rates provide insufficient correction.
@@
 Figure 2: Training dynamics across three embedding encoders on 2M multilingual samples.
-Qwen3-Embedding reaches 81.3% token accuracy at 72.5K steps with validation loss 1.32.
-All models show diminishing returns beyond 50K steps, suggesting architectural improvements rather than extended training as the path to further gains.
-5 Conclusion
-We presented embedding inversion via conditional masked diffusion, achieving 81.3% token accuracy across three embedding models with a 78M parameter decoder that requires no access to the target encoder.
-The progression from Vec2Text, which requires a compatible encoder architecture, to ALGEN, which needs alignment training, to Zero2Text, which queries the target API, to our work, which needs none of these, demonstrates that inversion attacks are becoming more accessible, not less.
-Production systems that cache or transmit embedding vectors under the assumption of irreversibility should treat them as sensitive data requiring protection equivalent to the original text.
-Current limitations include the restriction to 32-token sequences, whereas real documents are substantially longer.
-Scaling to longer sequences via hierarchical diffusion or sliding window approaches is a natural next step.
-Incorporating language model priors through classifier-free guidance could further improve reconstruction quality, particularly for low-frequency tokens.
-On the defense side, encrypted vector search and inversion-resistant embeddings through adversarial training or differential privacy on the embedding space remain important open problems for security-critical deployments.
+Qwen3-Embedding reaches 81.3% training token accuracy at 72.5K steps with validation loss 1.32.
+All models show plateauing behavior beyond 50K steps, indicating that further improvements may require architectural changes rather than extended training.
+5 Conclusion and Future Work
+We presented embedding inversion via conditional masked diffusion, demonstrating that parallel denoising can recover tokens from embedding vectors without access to the target encoder.
+The method operates in a strictly black-box setting: the embedding vector is the only input at inference time, with no correction loop through the encoder.
+This threat model is more restrictive than Vec2Text, which requires encoder access for iterative refinement, or ALGEN, which needs alignment training on encoder outputs.
+This architectural difference also explains the performance gap.
+Autoregressive methods with iterative correction benefit from a closed-loop feedback signal.
+Each refinement step re-embeds the current hypothesis and computes an explicit error signal against the target.
+Diffusion-based inversion lacks this mechanism.
+The embedding conditions the generation through AdaLN modulation but never verifies whether the output actually maps back to the target embedding.
+Errors compound without correction, which our ablation results confirm: dynamic masking schedules that provide implicit curriculum learning substantially outperform fixed schedules, partially compensating for the absence of explicit feedback.
+Two directions appear promising for closing this gap.
+Classifier-free guidance could inject stronger embedding signal during sampling without requiring encoder access at inference time.
+Alternatively, combining diffusion-based initialization with a lightweight correction step through a distilled or approximate encoder could recover much of the closed-loop advantage while maintaining broader applicability across encoder architectures.
+Scaling to longer sequences via hierarchical diffusion over sentence-level chunks is a natural next step.
+Appendix A Mask Ratio Ablation with mmBERT Decoder
+We conduct a separate ablation study using a different decoder architecture to investigate the effect of fixed mask ratios on training dynamics.
+Unlike the main experiments, which use an 8-layer transformer trained from scratch with the GPT-2 tokenizer, these experiments use a 22-layer transformer initialized from mmBERT pretrained weights with a 256K multilingual vocabulary.
+Five models were trained with fixed mask percentages of 10%, 20%, 40%, 80%, and 100%, sharing identical architecture (388M parameters), training configuration (batch size 380, learning rate $10^{-4}$, warmup 2000 steps), and pretrained initialization.
+Table 4: Mask ratio ablation with mmBERT decoder. All models early-stopped before 200K steps.
+Mask % Best Step Best Val Loss Final Train Acc Tokens (M)
+10% 43,500 5.9943 30.2% 3,413
+20% 42,500 5.9197 21.0% 3,284
+40% 38,500 5.8730 14.0% 3,174
+80% 35,500 5.7933 10.8% 2,955
+100% 30,500 5.8410 10.1% 2,590
+Higher mask ratios achieve lower validation loss, with 80% masking reaching 5.79 (Table 4).
+Training accuracy inversely correlates with mask ratio as expected.
+Higher mask ratios also converge faster, with 100% masking reaching its best validation loss at step 30,500 versus 43,500 for 10% masking.
+However, all fixed-mask models underperform configurations using the log-linear schedule from the main experiments, suggesting that dynamic masking provides curriculum learning benefits.
+(a) Training accuracy curves
+(b) Validation loss curves
+Figure 3: Training dynamics across fixed mask ratios with mmBERT decoder.
+All models plateau early, and higher training accuracy from lower mask ratios does not translate into better validation performance.